OpenStudy (anonymous):

write an equation of a line parallel to y=2/3x-5 in slope intercept form that passes through the point (-6, -1)

Nnesha (nnesha):

y = 2/3x - 5 do you know what is slope in this equation ??

OpenStudy (anonymous):

Do you know what slope int form is?

OpenStudy (anonymous):

@Nnesha the slope in the equation is 2/3

syed98 (syedmohammed98):

it's parallel so the gradient is same i:e 2/3 to fine the equation use the formula 

Nnesha (nnesha):

yes right so parallel lines should have same slope

Nnesha (nnesha):

slope of new equation is going to be 2/3 bec u have to writen equation of parallel line and parallel lines have same slope y = 2/3x + b now use that order pair where first values is x and 2nd one is y like this (x ,y) now replace x and y by their values then solve for b

Nnesha (nnesha):

write*

syed98 (syedmohammed98):

gradient/ slope ..same thing :)

OpenStudy (anonymous):

@Nnesha I don't really understand how to solve for b

Nnesha (nnesha):

first replace x and y by their values

syed98 (syedmohammed98):

:l

Nnesha (nnesha):

let me know how u plugged that into the equation so we can ssolve for b :)

OpenStudy (anonymous):

(-6, -1) = (x, y) ? @Nnesha

OpenStudy (anonymous):

is the answer y=2/3x-6 @Nnesha

Nnesha (nnesha):

nope y-intercept is not -6

Nnesha (nnesha):

 replace x by -6 and y by -1